# Toward Learning POMDPs Beyond Full-Rank Actions and State Observability

### v1

@@ Abstract
We are interested in enabling autonomous agents to learn and reason about systems with hidden states, such as furniture with hidden locking mechanisms. We cast this problem as learning the parameters of a discrete Partially Observable Markov Decision Process (POMDP). The agent begins with knowledge of the POMDP’s actions and observation spaces, but not its state space, transitions, or observation models. These properties must be constructed from action-observation sequences. Spectral approaches to learning models of partially observable domains, such as learning Predictive State Representations (PSRs), are known to directly estimate the number of hidden states. These methods cannot, however, yield direct estimates of transition and observation likelihoods, which are important for many downstream reasoning tasks. Other approaches leverage tensor decompositions to estimate transition and observation likelihoods but often assume full state observability and full-rank transition matrices for all actions. To relax these assumptions, we study how PSRs learn transition and observation matrices up to a similarity transform, which may be estimated via tensor methods. Our method learns observation matrices and transition matrices up to a partition of states, where the states in a single partition have the same observation distributions corresponding to actions whose transition matrices are full-rank. Our experiments suggest that these partition-level transition models learned by our method, with a sufficient amount of data, meets the performance of PSRs as models to be used by standard sampling-based POMDP solvers. Furthermore, the explicit observation and transition likelihoods can be leveraged to specify planner behavior after the model has been learned.
@@ 1 Introduction
When planning and acting in the real world, intelligent agents must learn and reason about hidden information. Of great inspiration to us is the work of baumOpeningLockbox2017, which shows that a real autonomous robot can infer a cabinet’s locking mechanism from a hypothesis set of possible mechanisms through interaction. We are interested in a more general problem where autonomous agents must learn, through interaction, the dynamics of a system with hidden states, without any knowledge of the system state and transitions beforehand. The agent should also compute explicit estimates of transition and observation likelihoods to support downstream operations that manipulate the model, such as the specification of tasks to direct agent behavior. Our problem is modeled as learning the parameters of a discrete Partially Observable Markov Decision Process (POMDP) from a sequence of actions and observations acquired through random exploration.
@@
One common approach to learning a representation of a probabilistic latent-variable models like a POMDP is to apply a spectral decomposition to a matrix that contains estimates of the joint likelihoods of the observable random variables (hsuSpectralAlgorithm2012; balleSpectralLearning2014). In particular, singular-value decompositions (SVD) give a way of estimating the number of hidden variables of the system by truncating the singular values under a certain threshold. For POMDPs, spectral methods may be applied to a Hankel matrix, which stores the joint likelihood between past and future observations conditioned on a sequence of past and future actions. The decomposition of this matrix can be used to derive a (linear) Predictive State Representation, which can be interpreted as an automaton with real-valued transition matrices (bootsClosingLearningplanning2011; balleSpectralLearning2014). The ‘state’ of the PSR is a sufficient statistic that can be used to predict the likelihood of future observations given a possible sequence of actions. This prediction capability allows PSRs to be used as black-box models for reinforcement learning (liuWhenPartially2022; zhanPACReinforcement2022), but since transition and observation likelihoods cannot be directly read from a PSR, these models are difficult to manipulate after they are learned.
@@
There are other POMDP-learning algorithms that yield estimates of observation and transition likelihoods, but under assumptions that ultimately restrict the class of POMDPs that can be learned. Approaches introduced by azizzadenesheliReinforcementLearning2016 and guoPACRL2016 utilize tensor decompositions to recover observation distributions for each action whose transition matrix is full-rank. To recover the transitions, however, these approaches must also make the assumption that for each action, the corresponding observation distribution must be unique for every state. Full-rank transitions are common when modeling many real-world POMDPs, especially when actions may ‘fail’ with some probability, causing the system state to self-transition. Many systems, however, do not have distinct observation distributions associated with every action, like the locking mechanisms of baumOpeningLockbox2017 or many standard POMDPs in the literature, like Tiger (kaelbling_planning_1998).
@@
We investigate the relationship between PSRs and tensor decomposition methods to learn a broader class of POMDPs than existing tensor methods. To connect the two approaches, we extend a result established by carlyleRealizationsStochastic1971a and balleSpectralLearning2014 that states that PSRs learn transitions and observation matrices up to an unknown basis. We then reformulate tensor decomposition methods to estimate the unknown basis to recover the original basis. Our modification of tensor decomposition methods for hidden state inference allows us to simultaneously leverage all observation distributions from all actions with full-rank transition methods all at once, rather than a per-action basis like previous approaches (azizzadenesheliReinforcementLearning2016; guoPACRL2016). Should the collection of observation distributions of all full-rank actions be unique for each state, like Tiger, we may recover the full POMDP. Should there exist states that share the same set of observation distribution when aggregated across actions, we learn transitions between partitions of states, where states in a single partition share the same observation distributions over all actions.
@@
Learning explicit transition and observation models is valuable because they enable model-based reasoning over environment dynamics. Whereas black-box PSRs only provide predictive likelihoods of observation sequences, access to explicit transition matrices and observation matrices allows for the specification of rewards after the model has been learned to direct planner behavior. Our experimental results suggest that our method can correctly learn partition-level transitions and observations and that these likelihoods are necessary to correctly direct agent behavior in POMDPs with very noisy observations.
@@ 2 Problem Setting
We assume that the ground truth system can be described as a discrete POMDP, represented by a 8-tuple $(\mathcal{S},\mathcal{T},\mathcal{A},\mathcal{O},\mathcal{Z},b_{0},R,\gamma)$. The set $\mathcal{S}=\{s^{1},s^{2},\dots\}$ is a discrete set of states, $\mathcal{A}$ is a discrete set of actions, and $\mathcal{O}=\{o^{1},o^{2},\dots\}$ is a discrete set of observations. $\mathcal{T}=\{T^{a}:a\in\mathcal{A}\}$ denotes a set of row-stochastic state transition matrices. The element $T^{a}_{ij}=\operatorname{\mathrm{P}}(s_{t+1}=s^{j}|s_{t}=s^{i},a_{t}=a)$ denotes the probability of transition to state $s^{j}$ from state $s^{i}$ after taking action $a$ at time $t$. The set $\mathcal{Z}=\{O^{ao}:(a,o)\in\mathcal{A}\times\mathcal{O}\}$ describes a collection of diagonal matrices, where $O^{ao}_{ii}=\operatorname{\mathrm{P}}(o_{t}=o|s_{t}=s^{i},a_{t}=a)$ denotes the emission probability of $o$ under action $a$ when leaving state $s^{i}$. The distribution $b_{0}\in\Delta(\mathcal{S})$ describes the distribution over the initial state. The constant $\gamma\in(0,1)$ is the reward discount factor.
The agent begins acting in a POMDP with access to the action and observation spaces $\mathcal{A}$ and $\mathcal{O}$. Under a uniform, memoryless random exploration policy $a_{t}\sim\operatorname{Unif}(\mathcal{A})$ for all $t\geq 1$, the agent collects a dataset $\mathcal{D}$, which is a long string of actions and observations $\mathcal{D}=(a_{1},o_{1},a_{2},o_{2},\dots)$. From this data, we wish the agent to estimate the number of hidden states $|\mathcal{S}|$, transition matrices $\mathcal{\hat{T}}=\{\hat{T}^{a}:a\in\mathcal{A}\}$, and observation matrices $\mathcal{\hat{Z}}=\{\hat{O}^{ao}:(a,o)\in\mathcal{A}\times\mathcal{O}\}$. We may also require the agent to learn a tabular reward $R$ function by including rewards as observations (izadiPointBasedPlanning2008). One way we evaluate the approach is by measuring the error of the estimated model parameters against those of the groundtruth POMDP. Another is by evaluating the performance agent behavior under a planning algorithm after the POMDP is inferred from $\mathcal{D}$. The last is by evaluating the behavior of a planner at a task designated by a user after the model has been learned.
@@ 3.1 Forward, Backward, and Hankel Matrices
To estimate systems of hidden state, a natural place to start is to form an array that expresses the joint likelihoods between the observable random variables. A Hankel matrix is an instance of these arrays that encodes the joint likelihoods of past and future action-observation trajectories. In this section, we derive the Hankel matrix given knowledge of the ground truth POMDP. Our construction starts with two intermediate factors, called the forward and backward matrix, which we will multiply together to form the Hankel matrix.
The forward matrix $\mathbf{Forw}$ is a $\infty\times{|\mathcal{S}|}$ matrix whose rows are indexed by histories of action-observation sequences. The row indices are determined by choosing an ordering that enumerates all possible history sequences. A sensible ordering is to enumerate the action-observation sequences of length one first, then of length two, etc. In practice, while there are an infinite number of action-observation sequences, enumerating sequences up to a certain length is sufficient for computation. If $\mathit{hist}=(a_{1},o_{1},\dots,a_{t},o_{t})$, then an entry $\mathbf{Forw}_{\mathit{hist},i}$ expresses the joint likelihood of observing the history and landing in state $s^{i}$. The entire row of $\mathit{hist}$ may be computed by ‘forward multiplying’ the POMDP matrices that correspond to the action and observations in $\mathit{hist}$
$\mathbf{Forw}_{\mathit{hist},i}$ $=\operatorname{\mathrm{P}}(o_{1},\dots,o_{t},s_{t+1}=s^{i}|a_{1},\dots,a_{t}),$ (1)
$\mathbf{Forw}_{\mathit{hist},:}$ $=b_{0}\cdot O^{a_{1}o_{1}}T^{a_{1}}\cdots O^{a_{n}o_{n}}T^{a_{n}}.$ (2)
The backward matrix $\mathbf{Back}$ is a $|\mathcal{S}|\times\infty$ matrix whose columns are indexed by ‘future’ action-observation sequences. Following the PSR literature (singhPredictiveState2004), we call these sequences tests. An ordering must be decided to determine the test indices in the same manner as the histories. Should $\mathit{test}=(a_{t+1},o_{t+1},\dots,a_{t+n},o_{t+n})$, then the entry $\mathbf{Back}_{i,\mathit{test}}$ expresses the likelihood of observing the test conditioned on beginning in state $s^{i}$. The entire column of $\mathit{test}$ can also be derived through forward multiplication of the POMDP matrices
$\mathbf{Back}_{i,\mathit{test}}$ $=\operatorname{\mathrm{P}}(o_{t+1},\dots,o_{t+n}|a_{t+1}\dots,a_{t+n},s_{t+1}=s^{i})$ (3)
$\mathbf{Back}_{:,\mathit{test}}$ $=O^{a_{t+1}o_{t+1}}T^{a_{t+1}}\cdots O^{a_{t+n}o_{t+n}}T^{o_{t+n}}\cdot\mathbf{1}$ (4)
where $\mathbf{1}$ is the vector where all entries are set to one.
@@
The Hankel matrix does not refer to the underlying hidden state of the POMDP and can be estimated from action-observation trajectories. If we had a long string of actions and observations $\mathcal{D}_{n}=(a_{1},o_{1},\dots,a_{n},o_{n})$, the matrix $\mathcal{H}$ could be estimated by the suffix-history approach, taking frequency counts of subsequences of increasing lengths (wolfeLearningPredictive2005; bootsClosingLearningplanning2011):
@@
where $\text{acts}(\mathit{hist}\oplus\mathit{test})$ is the action sequence associated with $\mathit{hist}\oplus\mathcal{D_{\mathrm{test}}}$ and $L=|\mathit{hist}\oplus\mathit{test}|<n$.
@@
This factorized construction is inspired by the construction of a related matrix, called the System Dynamics Matrix, by singhPredictiveState2004.111The entries of the System Dynamics Matrix (SDM) are the likelihood of a given test conditioned on a history. Each row of the Hankel matrix is the same as the SDM except scaled by a constant (the likelihood of the history that indexes the row; baconLearningPlanning2015). It is important to note that expressing the Hankel matrix as a factorization of $\mathbf{Forw}$ and $\mathbf{Back}$ represents the system under a memoryless policy where future actions are independent of previous observations. To correctly estimate the matrix via Eq. 6, the data must also be collected under a memoryless policy, such as the uniform exploration policy as introduced in Sec. 2 (bowlingLearningPredictive2006).
3.2 Transforming Predictive State Representations

### v2

@@ Abstract
We are interested in enabling autonomous agents to learn and reason about systems with hidden states, such as locking mechanisms. We cast this problem as learning the parameters of a discrete Partially Observable Markov Decision Process (POMDP). The agent begins with knowledge of the POMDP’s actions and observation spaces, but not its state space, transitions, or observation models. These properties must be constructed from a sequence of actions and observations. Spectral approaches to learning models of partially observable domains, such as Predictive State Representations (PSRs), learn representations of state that are sufficient to predict future outcomes. PSR models, however, do not have explicit transition and observation system models that can be used with different reward functions to solve different planning problems. Under a mild set of rankness assumptions on the products of transition and observation matrices, we show how PSRs learn POMDP matrices up to a similarity transform, and this transform may be estimated via tensor decomposition methods. Our method learns observation matrices and transition matrices up to a partition of states, where the states in a single partition have the same observation distributions corresponding to actions whose transition matrices are full-rank. Our experiments suggest that explicit observation and transition likelihoods can be leveraged to generate new plans for different goals and reward functions after the model has been learned. We also show that learning a POMDP beyond a partition of states is impossible from sequential data by constructing two POMDPs that agree on all observation distributions but differ in their transition dynamics.
@@ 1 Introduction
When planning and acting in the real world, intelligent agents must learn and reason about hidden information. Of great inspiration to us is the work of Baum et al. [baumOpeningLockbox2017], which shows that a real autonomous robot can infer a cabinet’s locking mechanism from a hypothesis set of possible mechanisms through interaction. We are interested in a more general problem where autonomous agents must learn, through interaction, the dynamics of a system with hidden states, without any knowledge of the system state and transitions beforehand. The agent should also compute explicit estimates of transition and observation likelihoods to support downstream operations that manipulate the model, such as task specification to direct agent behavior. Our problem is modeled as learning the parameters of a discrete Partially Observable Markov Decision Process (POMDP) from a sequence of actions and observations acquired through random exploration.
@@
One common approach to learning a representation of a probabilistic latent-variable model like a POMDP is to apply a spectral decomposition to a matrix that encodes correlations of the observable random variables [hsuSpectralAlgorithm2012, balleSpectralLearning2014]. For POMDPs, spectral methods may be applied to a Hankel matrix, which represents the correlation between past and future observations conditioned on a sequence of past and future actions. The decomposition of this matrix can be used to derive a (linear) Predictive State Representation [bootsClosingLearningplanning2011, balleSpectralLearning2014]. The ‘state’ of the PSR is a sufficient statistic that can be used to predict the likelihood of future observations given a possible sequence of actions. This prediction capability allows PSRs to be used as black-box models for reinforcement learning [liuWhenPartially2022, zhanPACReinforcement2022]; however, transition and observation likelihoods cannot be directly recovered from a PSR. This makes these models are difficult to manipulate, like changing the goal or reward function for planning. If the objective of the agent changes, the PSR must be relearned for the new task.
@@
There are other POMDP-learning algorithms that yield estimates of the full model, e.g. observation and transition likelihoods, but under assumptions that ultimately restrict the class of POMDPs that can be learned. Approaches introduced by Azizzadenesheli et al. [azizzadenesheliReinforcementLearning2016] and Guo et al. [guoPACRL2016] utilize tensor decompositions to recover observation distributions for each action whose transition matrix is full-rank. To recover the transitions, however, these approaches must also make the assumption that for each action, the corresponding diagonal observation matrices must be unique for every state, which implies every state has a unique observation distribution. While, full-rank transitions are common when modeling many real-world POMDPs, especially when actions may ‘fail’ with some probability, many real-world systems have aliased states, e.g. states that do not have distinct observation distributions associated with every action. Systems that fall in this class include the locking mechanisms of Baum et al. [baumOpeningLockbox2017] or many standard POMDPs in the literature, like Tiger [kaelbling_planning_1998].
@@
We investigate the relationship between PSRs and tensor decomposition methods to learn a broader class of POMDPs than existing tensor methods. A result established by Carlyle and Paz [carlyleRealizationsStochastic1971a] states that PSRs learn transitions and diagonal observation matrices up to an unknown basis. We then reformulate tensor decomposition methods to estimate the unknown basis to recover the original basis. Our modification of tensor decomposition methods for hidden state inference allows us to simultaneously leverage all observation distributions from all actions with full-rank transition methods all at once, rather than a per-action basis like previous approaches [azizzadenesheliReinforcementLearning2016, guoPACRL2016]. Should the collection of observation distributions of all full-rank actions be unique for each state, like Tiger, we may recover the full POMDP. Should there exist states that share the same set of observation distributions when aggregated across actions, we learn transitions between partitions of states, where states in a single partition share the same observation distributions over all actions, We also show that when restricted to sequential data, learning transition and observation up to observability partitions cannot be improved. We construct an example of two POMDPs whose dynamics differ between aliased states but yield the same distribution over all future observations under an arbitrary sequence of actions.
@@
Learning explicit transition and observation matrices is valuable because these models enable reasoning over environment dynamics. Whereas black-box PSRs only provide predictive likelihoods of observation sequences, access to explicit transition matrices and diagonal observation matrices allows for the specification of rewards after the model has been learned. Our experimental results suggest that our method can correctly learn partition-level transitions and observations and that these likelihoods are necessary to correctly direct agent behavior in POMDPs with very noisy observations.
@@ 2 Problem Setting
We assume that the ground truth system can be described as a discrete POMDP, a tuple $(\mathcal{S},\mathcal{T},\mathcal{A},\mathcal{O},\mathcal{Z},b_{0},R,\gamma)$. The set $\mathcal{S}=\{s^{1},s^{2},\dots\}$ is a discrete set of states, $\mathcal{A}$ is a discrete set of actions, and $\mathcal{O}=\{o^{1},o^{2},\dots\}$ is a discrete set of observations. $\mathcal{T}=\{T^{a}:a\in\mathcal{A}\}$ denotes a set of row-stochastic state transition matrices. The element $T^{a}_{ij}=\operatorname{\mathrm{P}}(s_{t+1}=s^{j}|s_{t}=s^{i},a_{t}=a)$ denotes the probability of transition to state $s^{j}$ from state $s^{i}$ after taking action $a$ at time $t$. The set $\mathcal{Z}=\{O^{ao}:(a,o)\in\mathcal{A}\times\mathcal{O}\}$ describes a collection of diagonal matrices, where $O^{ao}_{ii}=\operatorname{\mathrm{P}}(o_{t}=o|s_{t}=s^{i},a_{t}=a)$ denotes the emission probability of $o$ under action $a$ when leaving state $s^{i}$. The distribution $b_{0}\in\Delta(\mathcal{S})$ describes the distribution over the initial state. The constant $\gamma\in(0,1)$ is the reward discount factor.
The agent begins acting in a POMDP with access to the action and observation spaces $\mathcal{A}$ and $\mathcal{O}$. Under a uniform, memoryless random exploration policy $a_{t}\sim\operatorname{Unif}(\mathcal{A})$ for all $t\geq 1$, the agent collects a dataset $\mathcal{D}$, which is a long string of actions and observations $\mathcal{D}=(a_{1},o_{1},a_{2},o_{2},\dots)$. From this data, we wish the agent to estimate the number of hidden states $|\mathcal{S}|$, transition matrices $\mathcal{\hat{T}}=\{\hat{T}^{a}:a\in\mathcal{A}\}$, and diagonal observation matrices $\mathcal{\hat{Z}}=\{\hat{O}^{ao}:(a,o)\in\mathcal{A}\times\mathcal{O}\}$. We may also require the agent to learn a tabular reward $R$ function by including rewards as observations [izadiPointBasedPlanning2008]. We evaluate the approach is by measuring the error of the estimated model parameters against those of the ground-truth POMDP. Another is by evaluating the performance agent behavior under a planning algorithm after the POMDP is inferred from $\mathcal{D}$. The last is by evaluating the behavior of a planner at user-designated task after model learning.
For notational convenience, we will often need to refer to strings of actions and observations. We call a string $(a_{1},o_{1},\dots,a_{t},o_{t})$ that is observed by the agent in the past a history, often abbreviated as $\mathit{hist}$. A string $(a_{t},o_{t},\dots,a_{n},o_{n})$ the agent may observe in the future will be called a $\mathit{test}$ [littmanPredictiveRepresentations2001, singhPredictiveState2004]. To abbreviate the expression of the likelihood of observations conditioned on the actions, we write
$p(hist,test)$ $=\operatorname{\mathrm{P}}(o_{1},\dots,o_{t},o_{t+1},o_{n}|a_{1},\dots,a_{t},a_{t+1},\dots a_{n}),$
$p(hist,s_{t}=s^{i})$ $=\operatorname{\mathrm{P}}(o_{1},\dots,o_{t},s_{t}=s^{i}|a_{1},\dots,a_{t}),\text{ \ldots and so on.}$
Occasionally, it will be convenient to stack the diagonals of matrices in $\mathcal{Z}$ associated with the same action $a$ into $|\mathcal{S}|\times|\mathcal{O}|$ row-stochastic matrices $\text{Obs}^{a}$, where $\text{Obs}_{i,j}^{a}=\operatorname{\mathrm{P}}(o_{t}=o^{j}|s_{t}=s^{i},a_{t}=a)$. To distinguish between the two, we refer to the collection $\mathcal{Z}$ as diagonal observation matrices and the latter matrices $\{\text{Obs}^{a}:a\in\mathcal{A}\}$ as row-stochastic observation matrices.
@@ 3.1 Forward, Backward, and Hankel Matrices
To estimate systems of hidden state, a natural place to start is to form an array that expresses the correlation between the observable random variables. A Hankel matrix is an instance of these arrays that encodes the joint likelihoods of past and future action-observation trajectories. In this section, we derive the Hankel matrix given knowledge of the ground truth POMDP. Our construction starts with two intermediate factors, called the forward and backward matrix, which we will multiply together to form the Hankel matrix. [^1] This factorized construction is inspired by the construction of a related matrix, called the System Dynamics Matrix, by Singh et al. [singhPredictiveState2004]. The entries of the System Dynamics Matrix (SDM) are the likelihood of a given test conditioned on a history. Each row of the Hankel matrix is the same as the SDM except scaled by a constant, the likelihood of the history that indexes the row [baconLearningPlanning2015].
The forward matrix $\mathbf{Forw}$ is an infinite-by-$S$ matrix that expresses the joint likelihood of observing a history and the current state. After choosing a suitable ordering to map histories to indices (histories of length one, then histories of length two, and so on), we may write $\mathbf{Forw}_{hist,i}=p(hist,s_{t+1}=s^{i})$ for a given history $\mathit{hist}=(a_{1},o_{1},\dots,a_{t},o_{t})$. In terms of the POMDP matrices $\mathcal{T}$ and $\mathcal{Z}$, a row of the forward matrix can be expressed by right-multiplying the corresponding matrices in the order of the history:
$\mathbf{Forw}_{\mathit{hist},:}=b_{0}\cdot O^{a_{1},o_{1}}T^{a_{1}}\cdots O^{a_{t},o_{t}}T^{a_{t}}.$ (1)
The backward matrix $\mathbf{Back}$ is an $S$-by-infinite matrix that represents the conditional likelihood of obtaining the observations of a test after executing its actions, e.g. $\mathbf{Back}_{i,\mathit{test}}=p(test|s_{t+1}=s^{i})$. If the $\mathit{test}=(a_{t+1},o_{t+1},\dots,a_{n},o_{n})$, then a column of $\mathbf{Back}$ can also be computed similarly to the rows of $\mathbf{Forw}$:
$\mathbf{Back}_{:,\mathit{test}}=O^{a_{t+1},o_{t+1}}T^{a_{t+1}}\cdots O^{a_{n},o_{n}}T^{a_{n}}\cdot\mathbf{1},$ (2)
where $\mathbf{1}$ is a vector of one in all entries.
@@
The Hankel matrix does not refer to the underlying hidden state of the POMDP and can be estimated from action-observation trajectories. If we had a long string of actions and observations $\mathcal{D}_{n}=(a_{1},o_{1},\dots,a_{n},o_{n})$, the matrix $\mathcal{H}$ could be estimated by the suffix-history approach, taking frequency counts of subsequences of increasing lengths [wolfeLearningPredictive2005, bootsClosingLearningplanning2011]:
@@
where $\text{acts}(\mathit{hist}\oplus\mathit{test})$ is the action sequence associated with $\mathit{hist}\oplus\mathit{test}$, $\oplus$ is the concatenation operator, and $L=|\mathit{hist}\oplus\mathit{test}|<n$.
@@
It is important to note that expressing the Hankel matrix as a factorization of $\mathbf{Forw}$ and $\mathbf{Back}$ represents the system under a memoryless policy where future actions are independent of previous observations. To correctly estimate the matrix via Eq. 4, the data must also be collected under a memoryless policy. Using a uniformly random exploration policy admits this condition, but a larger class of non-memoryless exploration policies can be used for importance sampling [bowlingLearningPredictive2006].
3.2 Assumptions
[…]
Second, we also assume that $\mathbf{Forw}$, when limited to indices corresponding to one fewer than the maximum sequence length, has the same rank as the number of states (e.g. is full-rank), and that $\mathbf{Back}$ is also full-rank. This assumption is strictly weaker than prior work that assumes that all transition matrices $T^{a}$ and row-stochastic observation matrices $\text{Obs}^{a}$ are full rank for all actions [azizzadenesheliReinforcementLearning2016, guoPACRL2016]. The implication can be proven by showing that the products $\operatorname{diag}(b_{\pi})T^{a}\text{Obs}^{a}\operatorname{diag}(b_{\pi})^{-1}$ and $T^{a}\text{Obs}^{a}$ are submatrices of $\mathbf{Forw}$ and $\mathbf{Back}$, respectively. If $b_{\pi}$ is a stationary distribution, as implied by our ergodicity assumption, then every entry is positive, so $\mathbf{Forw}$ and $\mathbf{Back}$ that are full-rank. Many of the standard POMDPs in the literature, which are also included in our experiments (Sec. 5), are counterexamples for the converse of the implication.
Our assumptions have a few consequences on the estimated Hankel matrix. We only sample the starting state distribution $b_{0}$ at the start of the problem, so $b_{0}$ has little influence over the Hankel matrix. Instead, the Hankel matrix will take on the stationary distribution $b_{\pi}$ as the initial distribution instead. Furthermore, the rank of the resulting Hankel matrix will be equivalent to the number of states of the POMDPs in the restricted class that adhere to our assumptions, as opposed to the lower bound as is the case for general POMDPs [singhPredictiveState2004, huangMinimalRealization2016]. A proof of these claims is in Appendix A.1 of the supplemental material.
3.3 Constructing a Linear Predictive State Representation